What is the net income for the period year. The institution adjusts the books only at end of the year. The following is the unadjusted trial balance and some other info for December 31, 2019.

Accounts payable: 5,000
Accounts receivable: 4,000
Cash: 26,000
Common stock: 20,000
Dividends: 3,000
equipment: 16,000
Government contract revenue: 17,000
Prepaid rent: 4,000
Private sector revenue: 8,000
Retained earnings: 10,000
supplies: 2,000
Unearned revenue: 5,000
Wages expense: 10,000

1. Prepaid rent was recorded on November 21, when the company paid four months rent in advance for December 1st through march 31st.
2. A count of supplies on Dec 31 determined that 850 remained.
3. In early Dec, the company received 5,000 from the state of MA for services in advance. During the month, the institution earned 70% of the revenue.
4. Wages of 14,000 are paid at the end of every week. The current pay period is Dec 30th - Jan 5th.

After adjusting entries, what is the institution’s net income for the period year?
Answer #1

Adjusting entries:

No. General Journal Debit Credit 1 1,000 Rent expense Prepaid expense 1,000 2 1,150 Supplies expense Supplies 1,150 3 3,500 U


1. Prepaid rent for 4 months = 4,000
One month rent = 1,000

Adjustment of rent expense for the month of December = 1,000.
Therefore, rent expense is debited and prepaid rent is credited.

2. Supplies on hand at the beginning = 2,000
Supplies remaining on Dec 31 = 850
Supplies expense during the year = 2,000 - 850 = 1,150

Therefore, supplies expense is debited and supplies is credited.

3. Unearned revenue = 5,000
Unearned revenue earned during the month = 70%
Government contract revenue = 5,000 × 70% = 3,500

Therefore, unearned revenue is debited and Government contract revenue is credited.

4. Weekly wages = 14,000
Current pay period = Dec 30th - Jan 5th (7 days)
Daily wages = 14,000 / 7 = 2,000

Wages payable till Dec 31 (2 days) = 2,000 × 2 = 4,000

Therefore, wages expense is debited and wages payable is credited.


Income statement:

28,500 Income Statement Revenues: Government contract revenue 20,500 Private sector revenue 8,000 Total revenue Expenses: Wag


1. Government contract revenue = Existing balance + Adjusted balance = 17,000 + 3,500 = 20,500
2. Wages expense = Existing balance + Adjusted balance = 10,000 + 4,000 = 14,000
3. Rent expense for December = 1,000
4. Supplies expense = 2,000 - 850 = 1,150